OpenStudy (anonymous):

n(x+5)=ln(x-1)-ln(x+1) solve the logarithmic function algebraically. round the result to three decimal places.

OpenStudy (jamesj):

ok, what is ln a + ln b equal to?

OpenStudy (jamesj):

ln a + ln b = ...what?

OpenStudy (anonymous):

lnc?

OpenStudy (jamesj):

ln a + ln b = ln(ab) ln a - ln b = ... what?

OpenStudy (anonymous):

ln a-ln b= ln a divided by b?

OpenStudy (jamesj):

ln a - ln b = ln(a/b) yes Last identity before we solve your problem. If ln p = ln q what can we say about p and q?

OpenStudy (anonymous):

umm they can be multiplied or divide? idk

OpenStudy (jamesj):

If ln p = ln q then p = q

OpenStudy (jamesj):

Now, your problem: ln(x+5)=ln(x-1)-ln(x+1) Using the identity we wrote down above, what is ln(x-1)-ln(x+1) = ... ?

OpenStudy (anonymous):

lnx-1/lnx+1?

OpenStudy (jamesj):

No. ln a - ln b = ... what?

OpenStudy (anonymous):

ln(a/b)

OpenStudy (jamesj):

ln a - ln b = ln(a/b). hence ln(x-1)-ln(x+1) = ... what?

OpenStudy (anonymous):

ln(x-1/x+1)?

OpenStudy (jamesj):

Yes. Now \[ \ln(x+5)=\ln(x-1) - \ln(x+1) = \ln ( \frac{x-1}{x+1} ) \] which implies \[ x + 5 = \frac{x-1}{x+1} \] Now solve for x.

OpenStudy (anonymous):

ok thank yu so much